# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v2

@@ 4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. Firstly, we will verify the regularity of HPPP by an easy 1D example. Then, we compare the CP (14), PPP (8), GraRED-HP3 algorithms (see Algorithm 1), and GraRED-P3 (see Algorithm 2) for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms.
@@ 4.1 A toy example
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$ (See Figure 1).
Figure 1: Plot $f(x)+g(x)$.
Since for any $y\in\mathbb{R}$,
$\min\{(1+y)x-1,yx\}=\begin{cases}(1+y)x-1,&x<1\\ yx,&x\geq 1\end{cases},$
then $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y),y\in\mathbb{R}$, where
$\delta_{C}(y)=\begin{cases}0,&y\in C,\\ \infty,&y\notin C.\end{cases}$
Therefore, the corresponding saddle-point problem is:
$\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}xy+f(x)-g^{*}(y).$
We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fixed $x^{*}\geq 1$, then
$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0;\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
@@
For this toy example shown in Figure 2, we choose the same initial point $\mathbf{u}^{0}=(-6,6)$ or $(0,0)$, three different anchor points $\mathbf{a}=(12,10),(12,9),(12,8)$, the semi-definite $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, and the total itertation number $N=1000$. As shown in Figure 2, $\mathbf{u}^{k}$ generated by HPPP (5) can converge to the particular saddle point controlled by the anchor point. While the limit of $\mathbf{u}^{k}$ generated by PPP (8) may be uncertain, which is related with $\mathbf{u}^{0}$ and $\lambda_{k}$. For example, the sequence $\mathbf{u}^{k}$ oscillates around the limit $(1.8,0)$ shown in Figure 2(d).
Figure 2: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ in the $xOy$ Cartesian coordinate system.
As shown in Figure 3, $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}=\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be geometrically interpreted as $\mathcal{M}$-projection, i.e., the point $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ projection onto the line $x-y=x_{a}-y_{a}$ through the anchor point $\mathbf{a}=(x_{a},y_{a})$. Therefore, the $\mathcal{M}$-projection of $\mathbf{a}=(x_{a},y_{a})$ onto $\Omega$ is definitely solved by
$P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0;\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
@@ 4.2 Image deblurring
Firstly, we compared with classic TV regularization about CP [10], PPP [5] , and HPPP(see details in 3.3). For classic TV -$\ell^{2}$ regularization,
@@
where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator. In case $\mathbf{A}\mathbf{x}$ can be written as a convolution, i.e., $\mathbf{A}\mathbf{x}=\mathbf{k}\ast\mathbf{x}$, where $\mathbf{k}$ is the blurring convolution kernel. The primal-dual model of TV -$\ell^{2}$ deblurring is
(21) $\min_{\mathbf{u}\in\mathbb{R}^{n}}\max_{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{u},\mathrm{div}\mathbf{p}\rangle+\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2}-\delta_{P}(\mathbf{p}),$
where $f(\mathbf{u})=\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm defined as:
$\left\|\mathbf{p}\right\|_{\infty}=\max_{i,j}|\mathbf{p}_{i,j}|,|\mathbf{p}_{i,j}|=\sqrt{(\mathbf{p}_{i,j}^{1})^{2}+(\mathbf{p}_{i,j}^{2})^{2}}.$
We can calculate the proximal operator $(I+\tau\partial g^{*})^{-1}$ of the indicator $g^{*}(\mathbf{p})$, i.e.,
$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}.$
The resolvent operator for $f(\mathbf{x})$ can be computed by FFT.
$\mathbf{x}$ $=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})$
$=\arg\min_{\mathbf{x}}\frac{\|\mathbf{x}-\tilde{\mathbf{x}}\|}{2\tau}+\frac{\lambda}{2}\left\|\mathbf{k}*\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$
$=\mathcal{F}^{-1}\left(\frac{\tau\lambda\mathcal{F}(\mathbf{y})\mathcal{F}^{*}\left(\mathbf{k}\right)+\mathcal{F}(\tilde{\mathbf{u}})}{\tau\lambda\mathcal{F}\left(\mathbf{k}\right)^{2}+1}\right),$
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the FFT and inverse FFT, respectively.
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive WGN with noise level $0.01$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $\left\|K\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
@@
Secondly, we compared with RED and RED-PRO. The RED-PRO model uses the hybrid steepest descent method (HSD) [42, 43] to solve IR problems. Following [33, 14], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with an additive WGN with noise level $\sigma=\sqrt{2}$. The original RGB image is converted to the YCbCr image, PnP restoration algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images.
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3. The performance of three different methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularization is more powerful to regularize inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 4. To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 5, we plot their respective evolutions of PSNR values for iterations for the image House with 10 random initial points. The HPPP algorithm converges faster (less than 200 iterations) and achieves better PSNR values than CP and PPP algorithms. Once the anchor point is chosen, the proposed algorithm is more robust than CP and PPP algorithms for image deblurring with random initializations. To verify Corollary 3.7, we show the trend of convergence rate of the gap $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ between two successive iterations throughout the iterations with step size $\min\{\frac{2}{k},1\}(k\geq 1)$ in Figure 6.
@@
Figure 4: Deblurring result of Butterfly degraded with the Gaussian blur kernel with noise level $\sigma=0.01$.
Figure 5: Random initialization.
Figure 6: The convergence rate of $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ of GraRED-H3. The image Starfish is degraded by a Gaussian PSF with noise level $0.01$.
From the deblurring experiment Table 3 and Figure 7, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective.
Table 3: Recovery results are obtained by RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3 with TNRD [12].
@@
Figure 7: Deblurring result of Butterfly degraded by uniform kernel.
@@ 4.3 Image inpainting
In this section, we use the proposed algorithm to solve TV image inpainting problems and compare their numerical results with CP [10], PPP [5], and HPPP algorithms. The discrete image inpainting model is
@@
where $\left\|\nabla\mathbf{x}\right\|_{1}$ is the TV regularization, $\lambda,\beta$ are balance coefficients, $\odot$ indicates pointwise multiplication. Then the saddle-point problem is
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\max_{p\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{x},\mathrm{div}\mathbf{p}\rangle+\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}-\delta_{P}(\mathbf{p}),$
where $f(\mathbf{x})=\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm. Their resolvent operators of $f,g^{*}$ are
$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}$
and
$\mathbf{x}=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})=\frac{2\tau\mathbf{M}\odot\mathbf{y}+\tilde{\mathbf{x}}}{1+2\tau\mathbf{M}},$
the multiplication and division operators should be understood pointwise in the above formula.
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start their iterations with the degraded images. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters :
  • HPPP: $\tau=s=1/\left\|K\right\|=0.57$, anchor point $\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$, stepsize $\mu_{k}=\frac{1}{10(k+2)}$;
@@
In Table 4 and Table 5, we compared the numerical performance of classic algorithms with TV regularization. As we can see from both two tables, the proposed two algorithms outperform other algorithms. In Figure 8 and 9, we compare visualization results of House degraded by Bernoulli random mask and character mask, the proposed algorithms achieve better visual performance than TV regularization.
Moreover, we compare the recovery results about different anchors $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})$ with fixed $\mathbf{y}_{a}=\mathbf{0}$ and step size $\frac{1}{k+1}$. As shown in Figure 10, the anchor $\mathbf{x}_{a}=\mathbf{1}$ can achieve the best peroformance for random inpainting, which illustates the projection point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of $\mathbf{a}=(\mathbf{1},\mathbf{0})$ is closest to the true solution. Anchors selection is not difficult, GraRED-HP3 can achieve similar peroformance with other anchors.
Table 4: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, Bernoulli random missing).
@@
Table 5: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, character texture missing).
@@
Figure 8: Recovery results of House degraded by random mask.
@@
(h) GraRED-HP3 ($37.47$dB)
Figure 9: Recovery results of House degraded by character mask.
(a) Clean image
(b) Degraded image
(c) Anchor 1: 31.89dB
(d) Anchor 2: 31.95dB
(e) Anchor 3: 31.97dB
(f) Anchor 4: 31.85dB
Figure 10: Compared with different anchors for random inpainting, i.e., $\mathbf{x}_{a}=\mathbf{0},\mathbf{M}\odot\mathbf{x}_{0}+0.5(1-\mathbf{M}),\mathbf{1},\mathbf{x}_{0}$.

### v4

@@ 4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. Firstly, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and Restarted HPPP by other toy examples. Then, we compare the CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms. The source code is available at https://github.com/zsc15/HPPP.
@@ 4.1 A toy example
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, we plot $f(x)+g(x)$ in Figure 1.
[…]
Figure 1: The image of $f(x)+g(x)$.
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fixed $x^{*}\geq 1$, then
$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0,\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
@@
When $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, as shown in Figure 2, the minimization $\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}=\arg\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be interpreted as finding the $\mathcal{M}$-projection, namely, the projection of the anchor point $\mathbf{a}=(x_{a},y_{a})$ onto the line $x-y=x_{a}-y_{a}$. Therefore, the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\Omega$ is uniquely determined,
(21) $P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0,\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
$x-y-1\leq 0$$x$$y$$O$$S(2,0)$$R(3,1)$
@@
We set the total iteration number $N=1000$. As shown in Figure 3(a), with the anchor point fixed at $\mathbf{a}=(12,9)$, the sequence $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP converges to the same point $\mathbf{u}^{*}=(3,0)$, regardless of the initial point, which verifies Proposition 3.3 and (21). HPPP is also an implicitly regularized method. In contrast, Figure 3(b) shows that the sequence generated by PPP oscillates around the limit $(1.8,0)$. However, a quantitative mathematical characterization of the limit of the PPP sequence remains elusive.
Figure 3: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP or PPP.
4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41], for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
$T_{\theta}\mathbf{x}=\frac{1}{\gamma}\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\mathbf{x}.$
We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$) and Restarted HPPP ($q=50$). Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$. Figure 4 shows that HPPP and Restarted HPPP indeed provide accelerated residual decay.
Figure 4: Trajectories, distance to solution, two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
4.3 An infinite-dimensional example
We consider an infinite dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\cdots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
(22) $\mathbf{0}\in\mathcal{A}_{1}\mathbf{x}+\mathcal{A}_{2}\mathbf{x},\mathcal{A}_{1}=-\mathcal{S}-\mathbf{b},\mathcal{A}_{2}=I,$
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\cdots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\cdots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, then $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29]. Additionally, by solving the above inclusion problem (22) we obtain $\mathbf{x}^{*}=(1,0,\cdots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semi-definite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\cdots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
(e) Distance to solution
(f) Residual comparison
Figure 5: Comparison with PPP to solve an infinite dimensional problem in $\ell^{2}$ space.
@@ 4.4 Image deblurring
Firstly, we compared with CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
@@
where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator, and $\lambda,\beta$ are balance coefficients.
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $K=\left\|\nabla\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
@@
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3. The performance of these methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularizer is more powerful to regularize inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6. To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point. Once the anchor point is chosen, the proposed HPPP algorithm is robust to random perturbations of the initial points, which achieves stable recovery.
@@
Figure 6: Deblurring results of Butterfly degraded by the Gaussian blur kernel with noise level $\sigma=2.55$.
(a) Butterfly
(b) Parrots
Figure 7: Robustness comparison with random initialization. Test images are degraded by the Gaussian blurring kernel with the noise level $\sigma=2.55$.
Secondly, we compared with RED and RED-PRO. Following [47, 20], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted to the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and Restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, Restarted HPPP further boosts the performance of GraRED-HP3.
(a) Recovery peroformance
(b) Fixed point residual
Figure 8: Recovery peroformance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
Additionally, we trained an FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO. We also compared with state-of-the-art methods using advanced denoisers such as DRUNet ($D_{2}$[65] and GS denoiser ($D_{3}$[32, 33, 59]. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with TNRD [18]. The best two results are highlighted in red and blue colors, respectively.
@@
Table 4: Parameters on GraRED-HP3 and Restarted HPPP with advanced denoisers.
[…]
Restarted HPPP$(D_{2})$ $\tau=1.2,s=1/1.2,N=30,q=10$
[…]
Table 5: Compared with GraRED-P3, GraRED-HP3, and Restarted HPPP with TNRD for different tasks.
kernels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
noise level, scale $\sigma=5,\times 3$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$
GraRED-P3 27.49 29.25 30.92 28.19 33.53 33.73
GraRED-HP3 27.61 29.33 30.99 28.32 33.53 33.71
Restarted HPPP 27.66 29.36 31.00 28.34 33.53 33.73
Table 6: Compared with state-of-the-art methods. Test gray images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$. The best two results are highlighted in red and blue colors, respectively.
Pepper Craft Cameraman Couple Man House Starfish Butterfly Boat
RED [47] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
RED-PRO [20] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
DPIR [65] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
PnP-DRS [33] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
PnP-PGD [32] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
PnP-DYS [59] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
[…]
Restarted HPPP$(D_{2})$ 28.34 26.75 27.02 29.71 30.44 32.69 28.91 29.28 30.01
[…]
Restarted HPPP$(D_{3})$ 29.81 27.13 27.49 30.04 30.80 33.15 30.09 30.53 30.37
Figure 9: Deblurring results of Butterfly degraded by the uniform kernel.
@@ 4.5 Image inpainting
In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [16], PPP [9], and HPPP algorithms. The discrete image inpainting model is
@@
where $\mathbf{y}$ is the degraded image, $\mathbf{M}$ is a mask, and $\lambda,\beta$ are balance coefficients.
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start with the degraded image. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters:
  • HPPP: $\tau=s=1/\left\|K\right\|=0.57,\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a},\mu_{k}=\frac{1}{10(k+2)}$;
@@
In Table 7 and Table 8, we compared the numerical performance of classic algorithms with TV regularization. In Figure 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask, the proposed algorithms achieve better visual performance than TV regularization.
Table 7: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, Bernoulli random missing).
@@
Table 8: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, character texture missing).
@@
Figure 10: Recovery results of House degraded by the random mask.
@@
(p) GraRED-HP3 ($37.47$dB)
Figure 11: Recovery results of House degraded by the character mask.